# Superconductivity in the repulsive Hubbard model on different geometries induced by density-assisted hopping

### v1

@@ Appendix B Correlation functions
Bosonization predictions. In the C1S0 phase (Luther-Emery), bosonization predicts that the long-distance behavior of the density and pair correlation functions is governed by a power-law decay determined by the Luttinger liquid parameter $K_{\rho}$, i.e.,
@@
where $A_{s,d}$ and $B_{s,d,n}$ are non-universal constants and $n$ is the density. In this regime, the spin correlation function $C_{sz}$ decays exponentially. In the C1S1 phase (Luttinger liquid), the spin correlation function should also decay as a power law,
@@ Appendix E Scaling behavior of the momentum discontinuity $Z_{k_{F}}$
We analyze the finite-size scaling of the apparent momentum discontinuity $Z_{k_{F}}$ for our numerical result. In Fig. 6 we show its behavior for different values of $c_{n}$. We plot in panel (a) for selected values of $c_{n}$, $Z_{k_{F}}$ as a function of the system size in log-log scale. We can see that, as expected, all of them follow a power law $Al^{1-\gamma}$. In the Luther-Emery phase, we see the strongest system size effects. We present these results in panel (b), where $\gamma-1$ is plotted as a function of $c_{n}$. Inside the Luther-Emery phase ($c_{n}\gtrsim 0.23$), we see that the value of the exponent approximates to two ($\gamma\to 2$). For example, for $c_{n}=0.6$ the decaying exponent is $\gamma\simeq 1.925(5)$. On the other side of the transition, in the Luttinger liquid phase ($c_{n}\lesssim 0.23$), the decaying is much slower than on the attractive side, but still decaying as a power law. For example, for $c_{n}=0.0$, one can see a slow decaying behavior with $\gamma\simeq 1.048(2)$. In the vicinity of the transition, we see that this exponent becomes almost zero, showing almost no finite-size effects. As an example, we have $\gamma\simeq 1.00099(5)$ for $c_{n}=0.0$. This point is in the near vicinity of the effectively non-interacting point, and it shows a sharp stem function in its momentum distribution [shown in Fig. 2(b) in the main text].

### v2

@@ Appendix B Correlation functions
Bosonization predictions. We use the bosonization treatment of the effective Hamiltonian Eq. 2 in order to predict the properties of the system. The bosonized version of the Hamiltonian is given by
$H_{b}$ $=$ $\frac{1}{2\pi}\sum_{\nu=\rho,\sigma}\int dx\left[u_{\nu}K_{\nu}(\partial_{x}\theta_{\nu})^{2}+\frac{u_{\nu}}{K_{\nu}}(\partial_{x}\phi_{\nu})^{2}\right]$
$-\frac{2g_{1\perp}}{(2\pi\alpha)^{2}}\cos\sqrt{8}\phi_{\sigma}$
where $u_{\nu}$ are the mode velocities and $K_{\nu}$ the Luttinger parameters, $g_{1\perp}$ the backscattering, and $\phi_{\nu}$ and $\theta_{\nu}$ are, respectively, the bosonic field and its dual field (with $\nu=\rho,\sigma$ labeling the charge and spin sectors). Within this treatment one can predict the decay of the correlation functions in the different regimes. The charge mode $\phi_{\rho}$ is always gapless, but the renormalization group treatmentGiamarchi (2004) of the quantum sine-Gordon Hamiltonian describing the spin mode $\phi_{\sigma}$ identifies a gapped phase (C1S0) in which $e^{i\sqrt{2}n\phi_{\sigma}}$ has long range order and a gapless phase (C1S1) in which $g_{1\perp}\to 0$ and $K_{\sigma}\to 1$ so that $e^{i\sqrt{2}\phi_{\sigma}}$ has only quasi-long range order. In the C1S0 phase (Luther-Emery), bosonizationGiamarchi (2004) predicts that the long-distance behavior of the density and pair correlation functions is governed by a power-law decay determined by the Luttinger liquid parameter $K_{\rho}$, i.e.,
@@
where $A_{s,d}$ and $B_{s,d,n}$ are non-universal constants and $n$ is the density. In this regime, the spin correlation function $C_{sz}$ decays exponentially. In the C1S1 phase (Luttinger liquid), the spin correlation function should also decay as a power lawGiamarchi (2004), with logarithmic corrections coming from the renormalization flow
@@ Appendix E Scaling behavior of the momentum discontinuity $Z_{k_{F}}$
We analyze the finite-size scaling of the apparent momentum discontinuity $Z_{k_{F}}$ for our numerical result. In Fig. 6 we show its behavior for different values of $c_{n}$. We plot in panel (a) for selected values of $c_{n}$, $Z_{k_{F}}$ as a function of the system size in log-log scale. We can see that, as expected, all of them follow a power law $Al^{1-\gamma}$. In the Luther-Emery phase, we see the strongest system size effects. We present these results in panel (b), where $\gamma-1$ is plotted as a function of $c_{n}$. Inside the Luther-Emery phase ($c_{n}\gtrsim 0.23$), we see that the value of the exponent approximates to two ($\gamma\to 2$). For example, for $c_{n}=0.6$ the decaying exponent is $\gamma\simeq 1.925(5)$. On the other side of the transition, in the Luttinger liquid phase ($c_{n}\lesssim 0.23$), the decaying is much slower than on the attractive side, but still decaying as a power law. For example, for $c_{n}=0.0$, one can see a slow decaying behavior with $\gamma\simeq 1.048(2)$. In the vicinity of the transition, we see that this exponent becomes almost zero, showing almost no finite-size effects. As an example, we have $\gamma\simeq 1.00099(5)$ for $c_{n}=0.0$. This point is in the near vicinity of the effectively non-interacting point, and it shows a sharp step function in its momentum distribution [shown in Fig. 2(b) in the main text].
Appendix F Phase diagram for different parameters
For the main part of this work, we have chosen to show the data corresponding to a value of the Coulomb interaction $U=4t_{\parallel}$ and density $n=15/16=0.9375$, but the validity of the results is much wider than that. In Figs. 7 and 8, we present results for different interaction strengths ($U=8t_{\parallel}$, $12t_{\parallel}$, $16t_{\parallel}$, and $20t_{\parallel}$), and densities ($n=14/16=0.875$ and $13/16=0.8125$). The purple dashed lines in Fig. 7 correspond to the transition predicted by our effective model. Although the effective model is, in principle, valid for small values of $U$, the agreement seems to be very good up to large values of the interaction such as $U=12t_{\parallel}$ [Fig. 7(a-b)]. For larger values of the interaction, $U=16t_{\parallel}$ and $20t_{\parallel}$, the prediction of the critical $c_{n}$ is not very precise. Fermions occupy the antibonding band, and the interband terms in Eq. 1 must be taken into account. Despite this fact, in Fig. 8, we see that the key features discussed in the main text at the transition (the opening of the spin gap, the reduction of the central charge from $2$ to $1$, and a maximum on the discontinuity $Z_{k}$) remain the same even for very large values of the Coulomb interaction ($U=20t_{\parallel}$). Note that the maximum value of the discontinuity of the momentum distribution ($Z_{k}$) is reduced as a consequence of the reduced occupation of the lower band.
We have also probed the stability of the transition for different values of the density the results at $U=4t_{\parallel}$ for densities $n=14/16=0.875$ and $13/16=0.8125$ are presented in 7 and 8 in panels (e) and (f), respectively. In these cases, the transition predicted by the effective model agrees with the results found numerically, showing that the transition generically is stable in the studied parameter space.
Figure 7: The value of the central charge extracted from MPS simulations on 6464-rungs systems. The purple dashed lines correspond to the values of the transition predicted by the effective model. The corresponding simulation parameters are (a) $n=0.9375$, $U=8t_{\parallel}$, (b) $n=0.9375$, $U=12t_{\parallel}$, (c) $n=0.9375$, $U=16t_{\parallel}$, (d) $n=0.9375$, $U=20t_{\parallel}$, (e) $n=0.875$, $U=4t_{\parallel}$, and (f) $n=0.8125$, $U=4t_{\parallel}$.
Figure 8: Transition at $\tilde{t}_{\perp}=3t_{\parallel}$ for a 6464-rung ladder. (a) The gray and purple plots correspond to the central charge and the discontinuity of the momentum distribution $Z_{k_{F}}$, respectively. The orange plot corresponds to the inverse correlation length. The corresponding simulation parameters are (a) $n=0.9375$, $U=8t_{\parallel}$, (b) $n=0.9375$, $U=12t_{\parallel}$, (c) $n=0.9375$, $U=16t_{\parallel}$, (d) $n=0.9375$, $U=20t_{\parallel}$, (e) $n=0.875$, $U=4t_{\parallel}$, and (f) $n=0.8125$, $U=4t_{\parallel}$.